# Graph Convolutions Enrich the Self-Attention in Transformers!

### v2

@@ 4.3 Experiments on Automatic Speech Recognition
Experimental Setting.
We conduct automatic speech recognition (ASR) experiments on the LibriSpeech  [^1] http://www.openslr.org/12 dataset (Panayotov et al., 2015), which consists of audio recordings paired with their transcriptions. For implementation, we follow the recipes of SpeechBrain (Ravanelli et al., 2021) and describe the detailed settings in Appendix I.1. For our experiments, we use Branchformer (Peng et al., 2022), one of state-of-the-art models for speech recognition, as well as a pure Transformer.
@@ Results.
Table 4 compares word error rates (WERs) on LibriSpeech 100h and 960h. For LibriSpeech 100h, Transformer+GFSA achieves 10.30/25.30 on the test clean/other set, which is a 6.53% improvement over the Transformer model for the WER of test clean. For LibriSpeech 960h, Transformer+GFSA shows a WER result of 2.31 in test clean, a 4.55% improvement over Transformer. Figure 7 in Appendix I.2 depicts the learning curves of train loss and valid loss when using GFSA, showing the effectiveness of our proposed filter.
@@
Transformer + GFSA 71.5M 10.30 24.30 2.31 5.49
Branchformer (Peng et al., 2022) 109.8M 9.63 22.43 2.13 5.00
Branchformer + GFSA 109.8M 9.60 22.25 2.11 4.94
4.4 Experiments on Graph Transformers
Table 5: Experimental results and number of parameters on ZINC
Method #Params MAE
Graphormer 500K 0.1240
Graphormer + GFSA 500K 0.1189
Method #Params PCQM4M PCQM4Mv2
Train Validate Train Validate
Graphormer 48.3M 0.0535 0.1286 0.0250 0.0862
Graphormer + GFSA 48.3M 0.0312 0.1193 0.0249 0.0860
Experimental Setting.
To confirm the efficacy of GFSA on graph transformers, we conduct experiments on a commonly used graph-level prediction dataset (i.e., ZINC) from benchmarking-GNN leaderboards (Dwivedi et al., 2023) and the recent OGB-LSC quantum chemistry regression (i.e., PCQM4M-LSC) challenge (Hu et al., 2021), which is currently the biggest graph-level prediction dataset and contains more than 3.8M graphs in total. We choose Graphormer (Ying et al., 2021) as a backbone architecture and use its framework  [^2] https://github.com/microsoft/Graphormer for implementation. Mean absolute errors (MAEs) are reported for both regression tasks. For detailed experimental setting, see Appendix H.1.
@@
Tables 6 and 6 show that Graphormer with GFSA consistently outperforms the vanilla model across all datasets. As can be seen from the train MAE values of Table 6, using GFSA allows the model to learn more effectively during the training process. Notably, plugging GFSA improves the performance by a large margin, e.g., 7.20% relative validate MAE decline in PCQM4M. Due to space constraints, results with standard deviation are included in table 14 and  15 in Appendix H.2.
Table 7: Code defect accuracy for models with plugged in GFSA. The number in ($\uparrow$) indicates the improvement rate (%) over the base model.
Method Accuracy
RoBERTa (Liu et al., 2019) 62.88
RoBERTa + GFSA 64.39 ($\uparrow$ 2.40%)
CodeBERT (Feng et al., 2020) 64.31
CodeBERT + GFSA 64.49 ($\uparrow$ 0.12%)
PLBART (Ahmad et al., 2021) 62.63
PLBART + GFSA 62.96 ($\uparrow$ 0.52%)
CodeT5-small (Wang et al., 2021b) 63.25
CodeT5-small + GFSA 63.69 ($\uparrow$ 0.70%)
CodeT5-base (Wang et al., 2021b) 63.51
CodeT5-base + GFSA 64.75 ($\uparrow$ 1.95%)
4.5 Experiments on Code Classification
Experimental Setting.
We conduct a code defect detection task based on Devign dataset provided by (Zhou et al., 2019). We use RoBERTa (Liu et al., 2019), CodeBERT (Feng et al., 2020), PLBART (Ahmad et al., 2021), and CodeT5 (Wang et al., 2021b) as our backbone models. The detailed settings are in Appendix J.1.
Results.
@@
Table 7 shows the accuracy of all models; GFSA results better than the base models. The biggest improvement is 2.40% for RoBERTa. In the case of CodeT5-base, using GFSA shows an accuracy of 64.75, an improvement of 1.95% from 63.51. CodeT5-small+GFSA has only about 100 additional parameters compared to CodeT5-small with 60M parameters, and even more impressively, it surpasses the accuracy of CodeT5-base. In Appendix J.2, we include case studies for this task. We also report the results of code clone detection task in Appendix K. In Table 17, CodeT5-small+GFSA outperforms CodeT5-small with an F1 of 94.36.
4.6 Runtime Overheads
Our GFSA does not significantly increase training time. We summarize the runtime of base models vs. base models with GFSA in Appendix L. On GLUE tasks in natural language understanding, BERT with GFSA improves average performance from 82.51% to 83.58% (see Table. 1) with an overhead of less than 3 minutes in total training time. Considering the performance improvements, we think the increases in training time are negligible.
5 Final Remarks
@@
Our proposed graph filter-based self-attention (GFSA) mechanism achieves high performance with improvements on a variety of tasks in computer vision, natural language processing, graph pattern classification, speech recognition, and code classification. GFSA is a simple yet effective method that enriches the self-attention in Transformers with more diverse frequency information. This enables GFSA to address the oversmoothing problem and learn better latent representations for downstream tasks. However, our GFSA does not bring significant overheads in those Transformers’ empirical runtime complexity. One can use more complicated graph filters to enhance accuracy more, but our goal is to find a balance between accuracy enhancements and overheads in runtime complexity.
@@
We believe that GFSA is a promising new direction for improving Transformers. It is simple to implement and can be used in conjunction with other techniques to further improve the performance of Transformers. We hope that our work will inspire more research on graph filter-based self-attention and its applications.
Ethical Statements
In terms of the broader impact of this research on society, we do not see the very negative impacts that might be expected. However, this paper may have implications for the carbon footprint and accessibility of learning algorithms. The computations required for machine learning research are rapidly growing, resulting in a larger carbon footprint (Schwartz et al., 2020). Our study improves performance and increases runtime very slightly, but the runtime increase is not very significant. However, in future research, it will also be important to study and improve our GFSA by taking carbon footprints into account.
Reproducibility Statement
To ensure the reproducibility and completeness of this paper, we include the Appendix with 12 sections. Appendix A provides the complete proof of Theorem. 3.1. Appendix B contains the empirical study on the error related to Theorem. 3.1. Appendix C provides our PyTorch style pseudo code for our GFSA method. The pseudo code helps to implement our GFSA to any Transformers used a pure self-attention. All experiments in the paper are reproducible with additional implementation details provided in Appendices E to K.
@@
Appendix A Proof of Theorem. 3.1
@@
We begin by observing that the $L_{\infty}$ norm of $\bar{\bm{A}}^{2}$ is bounded by:
$\|\bar{\bm{A}}^{2}\|\leq\|\bar{\bm{A}}\|^{2}.$ (9)
By using the induction principle, it follows that for any positive integer $k$,
$\|\bar{\bm{A}}^{k}\|\leq\|\bar{\bm{A}}\|^{k}.$ (10)
Given that $\bar{\bm{A}}$ is a matrix normalized with softmax: i) all the elements of $\bar{\bm{A}}$ lie within [0, 1], and ii) the sum of all the elements in $\bar{\bm{A}}$ is equal to 1 (potentially less than 1 if an attention mask is used). Then we can calculate the upper bound of $L_{\infty}$ norm for $\bar{\bm{A}}$ as follows:
$\|\bar{\bm{A}}\|=\max_{i}\sum_{j}|\bar{\bm{A}}_{i,j}|\leq\max_{i}\mathbf{1}=1.$ (11)
@@
$E_{K}$ $=\|\bar{\bm{A}}^{K}-(\bar{\bm{A}}+(K-1)(\bar{\bm{A}}^{2}-\bar{\bm{A}}))\|$ (12)
$\leq\|\bar{\bm{A}}^{K}\|+\|\bar{\bm{A}}\|+(K-1)(\|\bar{\bm{A}}^{2}\|+\|\bar{\bm{A}}\|).$ (13)
From the prior expressions, we deduce that $\|\bar{\bm{A}}^{K}\|\leq\|\bar{\bm{A}}\|^{K}\leq 1^{K}=1$. Thus, the upper bound of the error term $E_{K}$ becomes as follows:
$E_{K}$ $\leq 1+1+(K-1)(1+1)=2K.$ (14)
Appendix B Empirical Error Bound
In this section, we empirically compare the theoretical error bounds in Theorem. 3.1 with the actual errors. Figure 3 shows the theoretical upper bound of $E_{K}$ and the actual $E_{K}$ for the attention of the last layer in a trained pure backbone. DeiT-S and Graphormer are trained on ImageNet-1k and ZINC, respectively, and BERT is finetuned on STS-B. It can be observed that for all models, the actual $E_{K}$ is always smaller than the theoretical error bound.
(a) DeiT-S
(b) BERT
(c) Graphormer
Figure 3: Theoretical and actual values of $E_{K}$
Appendix C Implementation of GFSA
The pseudo code of our GFSA is shown in Algorithm 1. For implememtation, $w_{0}$ and $w_{1}$ can be set as hyperparameters optionally.

### v5

@@ 5.5 Experiments on Automatic Speech Recognition
Setting.
We conduct automatic speech recognition (ASR) experiments on the LibriSpeech  [^3] http://www.openslr.org/12 dataset [55], which consists of audio recordings paired with their transcriptions. We use Branchformer [59] and a pure Transformer. For implementation, we follow the recipes of SpeechBrain [65] and the detailed settings are in Appendix N.1.
@@ Results.
Table 7 compares word error rates (WERs) on LibriSpeech 100h and 960h. For 100h, Transformer+GFSA achieves 10.30/25.30 on the test clean/other set, which is a 6.53% improvement over the Transformer for the WER of the test clean. For 960h, Transformer+GFSA shows a WER result of 2.31 in test clean, a 4.55% improvement over Transformer and Branchformer+GFSA achieves 2.31/5.49 with an LM on the test clean/other sets. Fig. 8 in Appendix N.2 depicts the learning curves of train loss and valid loss when using GFSA, showing the effectiveness of our proposed filter.
@@
+ GFSA 71.5M 10.30 24.30 2.31 5.49
Branchformer 109.8M 9.63 22.43 2.13 5.00
+ GFSA 109.8M 9.60 22.25 2.11 4.94
5.6 Experiments on Code Classification
Table 8: Results on code classification. The number in ($\uparrow$) indicates the improvement rate.
Method Accuracy
RoBERTa 62.88
+ GFSA 64.39 ($\uparrow$ 2.40%)
CodeBERT 64.31
+ GFSA 64.49 ($\uparrow$ 0.12%)
PLBART 62.63
+ GFSA 62.96 ($\uparrow$ 0.52%)
CodeT5-small 63.25
+ GFSA 63.69 ($\uparrow$ 0.70%)
CodeT5-base 63.51
+ GFSA 64.75 ($\uparrow$ 1.95%)
Setting.
We conduct a code defect detection task based on Devign dataset provided by Zhou et al. [105]. We use RoBERTa [44], CodeBERT [23], PLBART [2], and CodeT5 [84] as our backbone models. The detailed settings are in Appendix P.1.
@@
Table 8 shows the accuracy of all models; GFSA results better than the base models. The biggest improvement is 2.40% for RoBERTa. In the case of CodeT5-base, using GFSA shows an accuracy of 64.75, an improvement of 1.95% from 63.51. CodeT5-small+GFSA has only about 100 additional parameters compared to CodeT5-small with 60M parameters, and even more impressively, it surpasses the accuracy of CodeT5-base. The biggest improvement is 2.40% for RoBERTa. In Appendix P.2, we include case studies for this task. We also report the results of the code clone detection task in Appendix Q.
6 Discussion on Runtime Overheads
Limitation.
The introduction of our GFSA layer results in a slight increase in training and inference time. We report the runtimes when plugging GFSA in Appendices R and S. For GLUE benchmark, integrating GFSA into BERT enhances the average performance from 82.51% to 83.58% (see Table 1) with more overhead of less than 36 seconds per epoch based on average training time (see Table 23). Considering the improvements, the increases in training time are negligible.
GFSA in selected layers: a strategy to mitigate the limitation.
Figure 3: Effectiveness of our selective layer strategy on ImageNet-1k. This shows out strategy’s ability to maintain accuracy benefits while mitigating runtime increases.
As GFSA requires more calculation than the original self-attention, the runtime after using GFSA slightly increases. Our experiments initially applied GFSA across all Transformer layers (as discussed in Section 5); however, to reduce computational load, we propose a selective application strategy. For this purpose, GFSA is used only on even-numbered layers. In Tables 35 to 39 of Appendix T, the results show that this strategy effectively reduces runtime increases while preserving comparable performance to the full-layer GFSA integration. Notably, the selective application of GFSA cuts the per-epoch runtime increase by 26.90% relative to its full-layer application, with only a 7.39% increase in runtime per epoch compared to the backbone model in Table 39.
GFSA in linear Transformers.
Figure 4: Performance ($x$-axis), runtime ($y$-axis), and GPU usage (circle sizes) of various Transformers and integrated GFSA on Long-Range benchmark
Although GFSA requires additional computation for calculating $\bar{\bm{A}}^{2}$, we explore integrating GFSA with linear attention variants to maintatin efficiency and scalability. Recent approaches [36, 70] achieve linear complexity by reformulating softmax operations and reordering matrix multiplication in self-attention. We apply similar principles to compute second-order self-attention efficiently, enabling $\tilde{\bm{H}}_{\text{GFSA}}$ calculation with linear complexity with respect to sequence length. Fig. 4 shows the performance, runtime and GPU usage changes when applying our GFSA to Transformers with linear complexity. GFSA still improves performance compared to the backbone model, while the increase in time and GPU usage is minimal. Notably, when GFSA is applied to Efficient Attention [70], the performance is improved and the runtime is 11.82 times faster than when GFSA is applied to the vanilla self-attention. This shows that GFSA can be effectively implemented with linear complexity architectures while preserving its benefits and providing a solution for addressing computational concerns.
7 Conclusion
@@
Our proposed GFSA achieves high performance with improvements on a variety of tasks. GFSA is a simple yet effective method that enriches self-attention in Transformers with more diverse frequency information. This enables GFSA to address the oversmoothing problem and learn better latent representations for downstream tasks. However, our GFSA does not bring significant overheads in those Transformers’ empirical runtime complexity. One can use more complicated graph filters to enhance accuracy more, but our goal is to find a balance between accuracy enhancements and overheads in runtime complexity.
@@
We believe that GFSA suggests a promising new direction for improving Transformers. GFSA can be implement with simple way and used in conjunction with other techniques to further improve the performance of Transformers. Considering the ongoing advancements in large language models, such as GPT-4 [1] and LLaMA [76], we hope that our approach may offer new insights for enhancing their performance and efficiency.
Acknowledgements
N. Park was partly supported by the Korea Advanced Institute of Science and Technology (KAIST) grant funded by the Korea government (MSIT) (No. G04240001, Physics-inspired Deep Learning, 10%), Institute for Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (No. RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University), 20%; No. RS-2024-00457882, AI Research Hub Project, 50%), and Samsung Electronics Co., Ltd. (No. G01240136, KAIST Semiconductor Research Fund (2nd), 10%). K. Lee acknowledges support from the U.S. National Science Foundation under grant IIS 2338909. Dr. Trask acknowledges funding under the Department of Energy under the Mathematical Multifaceted Integrated Capability Centers program.
@@ Acknowledgements
Appendix A Reproducibility Statement
To ensure the reproducibility and completeness of this paper, we include the Appendix with 12 sections. Appendix I provides our PyTorch-style pseudo code for our GFSA method. The pseudo code helps to implement our GFSA to any Transformers used a pure self-attention. All experiments in the paper are reproducible with additional implementation details provided in Appendices K to Q.
Appendix B Broader Impact
In terms of the broader impact of this research on society, we do not see the very negative impacts that might be expected. However, this paper may have implications for the carbon footprint and accessibility of learning algorithms. The computations required for machine learning research are rapidly growing, resulting in a larger carbon footprint [69]. Our study improves performance and increases runtime very slightly, but the runtime increase is not very significant. However, in future research, it will also be important to study and improve our GFSA by taking carbon footprints into account.
Appendix C Oversmoothing and Additional Visualizations
In Fig. 2, we show the visualizations of oversmoothing characteristics in DeiT. We also provide visualizations in other domains. We show the filter response, cosine similarity, and singular value of BERT finetuned on STS-B dataset of GLUE tasks in Fig. 5 and Graphormer finetuned on ZINC dataset in Fig. 6.
To characterize self-attention, we first analyze the filter response of self-attention in the frequency domain. We follow the method used by Wang et al. [80] for spectral visualization of the self-attention matrix. As shown in Fig. 2 (a), DeiT has a near-zero magnitude for the high frequencies, which is characteristic of a low-frequency filter and is likely to result in oversmoothing when applied multiple times.
We follow the calculation method of Guo et al. [28] for cosine similarity. As shown in Fig. 2 (b), the higher similarity as the layers of the model get deeper is related to the oversmoothing problem. To further analyze this issue, we also consider the dimensionality collapse in Transformer-based models. We plot the singular value distribution of the feature in the last block. As shown in Fig. 2 (c), insignificant, near-zero values dominate the feature distribution. As layers get deeper, the similarity of features increases and dimensional collapse occurs. The oversmoothing problem is the same in BERT and Graphormer, as shown in Fig. 5 and Fig. 6.
(a) Filter reponse
(b) Cosine Similarity
(c) Singular Value
(a) Filter reponse
(b) Cosine Similarity
(c) Singular Value
Appendix D Analysis of Frequency Responses with Visualizations
We analyze the frequency responses, which represent the impact of learned coefficients, for all 12 layers of BERTBASE with and without GFSA. From Fig. 7, our analysis reveals that GFSA learns various filter types between layers. In early layers, we observe a tendency towards low-pass filtering, with prominent peaks at low frequencies. This aligns with the need for broader feature extraction in initial layers. The middle layers show a mix of low-pass and high-pass characteristics, with more complex frequency responses. This suggests GFSA is learning to balance between feature extraction and refinement. In deep layers, there is a noticeable shift towards higher frequency responses, indicating a move towards high-pass filtering. This shift supports our claim that GFSA can mitigate oversmoothing in deeper layers. BERTBASE+GFSA shows a consistently higher magnitude response at higher frequencies, especially in deeper layers, compared to vanilla BERT. In other word, vanilla self-attention works primarily as a low-pass filter, while GFSA utilizes a wider range of frequencies.
Figure 7: Visualization of the frequency responses for all 12 layers of BERT trained on STS-B dataset. The top-left figure corresponds to the first layer, and the bottom-right figure corresponds to the last layer.
Appendix E Frequency Analyses in the Singular Value Domain
Graph signal processing (GSP) [67, 68] can be understood as a generalized concept of DSP — in other words, DSP is a special case of GSP where a line graph with $n$ nodes is used and therefore, the graph Fourier transform (GFT) of the line graph is identical to the discrete Fourier transform.
However, one can use the singular value decomposition, when the GSO is not diagonalizable but symmetrically normalized, instead of the eigendecomposition [49]. Both the singular value decomposition and the eigendecomposition project the original signal onto a set of basis, but they use different basis sets. In the singular value decomposition, we sort the set of basis in ascending order of their eigenvalues, and perform frequency domain-like analyses [106, 49].
Since the self-attention matrix’s row-wise sum is always 1, the following is the case: $\bar{\bm{A}}=\bm{D}^{-1}\bm{A}=\frac{1}{n}\bm{A},$ where $n$ is the number of tokens. Maskey et al. [49] define the following symmetrically normalized adjacency matrix (SNA): $\bm{D}_{in}^{-1/2}\bm{A}\bm{D}_{out}^{-1/2}.$ Since the degree of every node is $n$ in the self-attention matrix, the following is the case: $\bm{D}_{in}^{-1/2}\bm{A}\bm{D}_{out}^{-1/2}=\bm{D}^{-1/2}\bm{A}\bm{D}^{-1/2}=\frac{1}{\sqrt{n}}\bm{A}\frac{1}{\sqrt{n}}=\frac{1}{n}\bm{A}=\bar{\bm{A}}.$ Therefore, the self-attention matrix is a special case of SNAs.
Appendix F Matrix Polynomial vs. Graph Fourier Transform
There are two paradigms of implementing graph filters: i) matrix polynomial, which does not require diagonalizability, and ii) graph Fourier transform, which uses the eigendecomposition for diagonalizable adjacency matrices or uses the Jordan decomposition or the singular value decomposition for non-diagonalizable adjacency matrices.
Those two paradigms have their own weaknesses: i) the matrix polynomial approach requires explicit matrix multiplications, and ii) the graph Fourier transform approach requires expansive spectral decompositions. The matrix polynomial is preferred when there are not many matrix multiplications. Otherwise, the graph Fourier transform approach may be better since the matrix multiplication can be simplified after the decomposition.
Appendix G Proof of Theorem 3.1
Theorem 3.1 (Filter characteristics based on coefficient values).
Let $\bar{\bm{A}}$ be a self-attention matrix interpreted as a graph with connected components. Consider the polynomial graph filter defined by $\sum_{k=0}^{K}w_{k}\bar{\bm{A}}^{k}$, where $w_{2},w_{3},\ldots,w_{K-1}=0$ and only $w_{0}$, $w_{1}$, and $w_{K}$ are non-zero. If the coefficients $w_{k}$ for $k=0,1,K$ are positive and their sum is 1, then the polynomial filter acts as a low-pass filter, attenuating high-frequency components and promoting smoothness across the graph. Conversely, if $w_{k}=(-\alpha)^{k}$ for $k=0,1,K$ and $\alpha\in(0,1)$ with sufficient large $K$, the polynomial filter exhibits high-pass filter behavior.
Note that without filtering, the singular value ratio is ${\left|\sigma_{i}^{0}\right|}/{\left|\sigma_{1}^{0}\right|}=1$. In the case where $\left|g(\sigma_{i})/g(\sigma_{1})\right|<1$ $\forall i\geq 2$, it implies that after applying the graph filter $g$, the lowest frequency component further dominates, indicating that the graph filter acts as a low-pass filter. Conversely, in the case where $\left|g(\sigma_{i})/g(\sigma_{1})\right|>1$ $\forall i\geq 2$, it implies that after applying the graph filter $g$, the lowest frequency component $\sigma_{i}$ no longer dominates, indicating that the graph filter acts as a high-pass filter.
@@ Appendix E Frequency Analyses in the Singular Value Domain
We prove the low-pass filter result. For the case where $w_{0}$, $w_{1}$, and $w_{K}$ are positive and their sum is $1$, we show that
$\left|g(\sigma_{1})\right|=\left|w_{0}+w_{1}+w_{K}\right|=1$ (12)
Hence, proving Theorem Theorem 3.1 is equivalent to show $\left|g(\sigma_{i})\right|<1$.
$\left|g(\sigma_{i})\right|=\left|w_{0}+w_{1}\sigma_{i}+w_{K}(\sigma_{i}+(K-1)(\sigma_{i}^{2}-\sigma_{i}))\right|<\left|w_{0}+w_{1}\sigma_{i}+w_{K}\sigma_{i}\right|=\sigma_{i}<1$ (13)
since $\sigma_{i}+(K-1)(\sigma_{i}^{2}-\sigma_{i})=\sigma_{i}((K-1)\sigma_{i}-(K-2))<\sigma_{i}((K-1)-(K-2))=\sigma_{i}$
For the high-pass filter result, when $w_{k}=(-\alpha)^{k}/(k+1)$ where $k=0,1,K$ and $\alpha\in(0,1)$, then we show that when $w_{0}=1,w_{1}=-\alpha/2$,
$\left|\frac{\lim_{K\rightarrow\infty}g(\sigma_{i})}{\lim_{K\rightarrow\infty}g(\sigma_{1})}\right|$ $=\left|\frac{\lim_{K\rightarrow\infty}w_{0}+w_{1}\sigma_{i}+w_{K}(\sigma_{i}+(K-1)(\sigma_{i}^{2}-\sigma_{i}))}{\lim_{K\rightarrow\infty}w_{0}+w_{1}+w_{K}(1+(K-1)(1-1))}\right|$ (14)
$=\left|\frac{\lim_{K\rightarrow\infty}1-\frac{\alpha}{2}\sigma_{i}+\frac{(-\alpha)^{K}}{(K+1)}(\sigma_{i}+(K-1)(\sigma_{i}^{2}-\sigma_{i}))}{\lim_{K\rightarrow\infty}1-\frac{\alpha}{2}+\frac{(-\alpha)^{K}}{(K+1)}}\right|$ (15)
$=\left|\frac{1-\frac{\alpha}{2}\sigma_{i}+(\lim_{K\rightarrow\infty}\frac{(-\alpha)^{K}}{(K+1)}(\sigma_{i}+(K-1)(\sigma_{i}^{2}-\sigma_{i})))}{1-\frac{\alpha}{2}}\right|$ (16)
$=\left|\frac{1-\frac{\alpha}{2}\sigma_{i}}{1-\frac{\alpha}{2}}\right|>1$ (17)
since
$\lim_{K\rightarrow\infty}\frac{(-\alpha)^{K}}{(K+1)}(\sigma_{i}+(K-1)(\sigma_{i}^{2}-\sigma_{i}))$ $=\lim_{K\rightarrow\infty}\frac{(-\alpha)^{K}}{(K+1)}(K-1)(\sigma_{i}^{2}-\sigma_{i})$ (18)
$=\lim_{K\rightarrow\infty}(-\alpha)^{K}\frac{(K+1)}{(K-1)}(\sigma_{i}^{2}-\sigma_{i})=0$ (19)
It shows that the graph filter with $w_{k}=(-\alpha)^{k}/(k+1)$ for $k=1,2,K$ emphasizes high-frequency components and acts as a high-pass filter.
This proof supports that the behavior of the polynomial filter as either a low-pass or high-pass filter directly depends on the sign and values of the coefficients, as specified in Theorem 3.1. ∎
Appendix H Proof of Theorem 4.1
Proof.
The Frobeinus norm of the self-attention is directly related to how far the softmax probabilities are from being uniform. For any matrix $\bm{M}\in\mathbb{R}^{m\times n}$, we have
$\|\text{softmax}(\bm{M})\|_{F}=\sqrt{\frac{m+\sum_{i=1}^{m}d_{\chi^{2}}(S_{i},U_{n})}{n}},$ (20)
where $S_{i}$ is the $i$-th row of $\text{softmax}(M)$, $U_{n}$ is the uniform distribution over $n$ elements, and $d_{\chi^{2}}(p,q)=\sum_{i}q_{i}(p_{i}/q_{i}-1)^{2}$ is the $\chi^{2}$-divergence between $p$ and $q$ [14]. The Frobenius norm is maximized when the whole mass of the probabilities is on one element, which is a case for $d_{\chi^{2}}(S_{i},U_{n})=n-1$ and $\|\text{softmax}(\bm{M})\|_{F}=\sqrt{m}$. Therefore, we can calculate the upper bound of Frobenius norm for $\bar{\bm{A}}$ as follows:
$\|\bar{\bm{A}}\|_{F}\leq\sqrt{n}.$ (21)
Note that $\bar{\bm{A}}\in\mathbb{R}^{n\times n}$ is a right-stochastic matrix normalized with row-wise softmax: i) all the elements of $\bar{\bm{A}}$ lie within [0, 1], and ii) the row-sum in $\bar{\bm{A}}$ is equal to 1. Since the self-attention matrix is a right-stochastic matrix, the power of the self-attention is also a right-stochastic matrix. Therefore, Eq. (21) is also hold for $\bm{\bar{A}}^{K}$ as follows:
$\|\bar{\bm{A}}^{K}\|_{F}=\sqrt{\sum_{i,j}\bar{\bm{A}}_{i,j}^{K}}\leq\sqrt{\sum_{i,j}\bar{\bm{A}}_{i,j}}=\|\bar{\bm{A}}\|_{F}\leq\sqrt{n}.$ (22)
@@
$E_{K}$ $=\|\bar{\bm{A}}^{K}-(\bar{\bm{A}}+(K-1)(\bar{\bm{A}}^{2}-\bar{\bm{A}}))\|_{F}$ (23)
$\leq\|\bar{\bm{A}}^{K}\|_{F}+\|\bar{\bm{A}}\|_{F}+(K-1)(\|\bar{\bm{A}}^{2}\|_{F}+\|\bar{\bm{A}}\|_{F})$ (24)
$\leq\sqrt{n}+\sqrt{n}+(K-1)(\sqrt{n}+\sqrt{n})=2\sqrt{n}K.$ (25)
Appendix I Implementation of GFSA
The pseudo code of our GFSA is shown in Algorithm 1. For implementation, $w_{0}$ and $w_{1}$ can be set as hyperparameters optionally.